OpenStudy (anonymous):

At the start of an experiment, 2x10^4 bacteria are present in a colony. Eight hours later, the population is 3x10^4. a.) Determine the growth constant k b.) What was the population two hours after the start of the experiment? c.)How long will it take for the population to triple?

OpenStudy (anonymous):

@wolf1728 Could you help me with this?

OpenStudy (aum):

Population \(P = A*e^{kt}\) when t = 0, P = 2x10^4. 2x10^4 = A * e^0. A = 2x10^4 \(P = 2\times10^4 * e^{kt} \) when t = 8, P = 3 x 10^4. Substitute and solve k which is the growth constant.

OpenStudy (anonymous):

oh okay makes sense

OpenStudy (anonymous):

@aum now for b.) i just useall that information into N(t)=Noe^kt

OpenStudy (anonymous):

correct?

OpenStudy (aum):

Yes. Once you find k from part a, you have the complete equation for P as a function of time t. For part b, all you have to do is set t = 2 and solve for P.

OpenStudy (wolf1728):

Okay the population goes from 20,000 to 30,000 in 8 hours Therefore it increases by a factor of 1.5 every 8 hours So the population increases by this equation: 20,000*1.5^(hours/8) zero hours 20,000*1.5^(0/8) = 20,000 two hours 20,000*1.5^(2/8) = 22,134

OpenStudy (anonymous):

ok

OpenStudy (wolf1728):

Time to triple is: 20,000*1.5^(X/8) = 60,000 1.5^(X/8) = 3 taking logs of both sides (X/8) *log(1.5) = log(3) (X/8) = log(3) / log(1.5) (X/8) = 2.7095112914 X = 21.6760903308 hours ... and it's just that simple :-) LOL

OpenStudy (wolf1728):

aum was quite helpful and used all that 'e' and natural log stuff I prefer more traditional (common log) methods